OpenStudy (anonymous):

If w varies directly as 2x - 1, and w = 9 when x = 2, find x when w = 15.

OpenStudy (anonymous):

jk im an idiot x = 3 doesn't it.

OpenStudy (kropot72):

You are correct.

OpenStudy (anonymous):

So then. wait. @kropot72 what is x when w is 15...

OpenStudy (kropot72):

w = k(2x - 1) where k is a constant of proportionality. 9 = k(4 - 1) = 3k k = 3 w = 3(2x - 1) 15 = 3(2x - 1)..............(1) Just solve equation (1) to find the value of x when w = 15

OpenStudy (anonymous):

@kropot72 you're actually perfect and I love you

OpenStudy (kropot72):

You're welcome :)